Two people are riding an elevator that can stop at floors two, three, …., and N. What is the probability that they will get off at different floors?
Answer #1

Number of ways they can get off at different floors = (N - 1)(N - 2)

Total number of ways = (N - 1)(N - 1)

Hence,

P(They will get off at different floors)

= (N - 2) / (N - 1)